OpenStudy (anonymous):

how do i solve this exponential equation... e^2x-4(e^x)=0

OpenStudy (anonymous):

e^(2x)-4(e^x)=0

jimthompson5910 (jim_thompson5910):

Let z = e^x, so z^2 = e^(2x) So your goal is to solve z^2 - 4z = 0 Then use the solutions in terms of 'z' to find the solutions in terms of 'x'

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

oh sorry, i understood :) thanks

jimthompson5910 (jim_thompson5910):

That's great. What did you get for the answers?

OpenStudy (anonymous):

i got x = ln4

jimthompson5910 (jim_thompson5910):

very nice, I'm getting that as well

OpenStudy (anonymous):

awesome